Grade 12th PassPhysical Chemistry

A sample of wustite FexO contains one Fe(III) for every three Fe(II). The value of x is

sachin kumar
8 Years agoGrade 12th Pass

3 Answers

Shreyash
ApprovedApproved Tutor Answer8 Years ago
According to the question, for x mole of Fe, x/4 mole Fe3+ and 3x/4 mole of Fe2+ are present in the sample.Balancing the charge, (x/4)*3+(3x/4)*2=2*1We get x=8/9.So the compound is Fe8O9 or Fe8/9O.Thanks for reading
Sandeep Bhamoo
8 Years ago
Let amount of Fe(+2)=3a. And, Fe(+3)=a. Then ,total (+)charge = total (-) charge-->. 3(a) +2(3a)=2. --> a = 2÷9 --> since total Fe = 3a + a = 4a --> Fe = 8÷9= .888
Sakshi
5 Years ago
The Gibbs energy for the decomposition al203 at 500 degrees celsius
2 by 3 a Al to O3 give reason for y3 a l + O2 the potential difference needed for electrolytic reduction of Al to O3 at 500 degrees celsius